OpenStudy (anonymous):

3/4-1/3x=1/4, solve for x? WHAT!?!?! Help, while explaining how to do it. please. :)

OpenStudy (anonymous):

subtract 3/4 from both sides. Then multiply both sides by -3

OpenStudy (anonymous):

polpak wht did u deleted my anser

OpenStudy (anonymous):

so that would mean, -1/3x = -1/2?

OpenStudy (anonymous):

Right so far.

jimthompson5910 (jim_thompson5910):

\[\frac{3}{4}-\frac{1}{3}x=\frac{1}{4}\] \[9-4x=3\] \[-4x=3-9\] \[-4x=-6\] \[x=\frac{-6}{-4}\] \[x=\frac{3}{2}\] Note: in the second step, I multiplied EVERY term by the LCD 12 to clear out the fractions.

OpenStudy (anonymous):

I deleted it because it was wrong, and because it wasn't what he asked for Hashir.

OpenStudy (anonymous):

Don't post answers only, and be sure that when you're explaining something you actually understand the question.

OpenStudy (anonymous):

that ws all giberish to me..

OpenStudy (anonymous):

Matt, if you go from what you have and multiply both sides by -3 you should get the answer jim got

OpenStudy (anonymous):

but u doesnot have to delete the answer because u doesnot have the rights too

OpenStudy (anonymous):

I do actually

OpenStudy (anonymous):

okay, got it.

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

Basically I cleared out the fractions first which gave me a simple linear equation. I then solved for x as I normally would with any normal equation.

OpenStudy (anonymous):

Yep, perfectly legit way to do it.